OpenStudy (anonymous):

help me in physics 2

OpenStudy (anonymous):



OpenStudy (anonymous):

A HOUSE USES the following appliances, aircon, (2hp0), ref (600 watts), rice cooker (1000 watts), sdvd (400 watts), stereo 50 watts, fla iron 500 watts, lamps 500 watts. if all the appliances are used daily withn an average of 3 hrs. how much will be paid in one month (30 days) assume that the electricity cost at an average of php 8.75/kwh

OpenStudy (anonymous):

what does that say, 2 hp , that means 2 horsepower

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

2 hp = 1 491.39974 watts according to google https://www.google.com/search?q=2+hp+%3D+watts&ie=utf-8&oe=utf-8

OpenStudy (anonymous):

im not sure what php stands for

OpenStudy (anonymous):

php stands pfor philippine peso

OpenStudy (anonymous):

1491.3 + 600 + 1000 + 400 + 50 + 500 + 500 watts = 13541.3 watts = 13.5413 kW 13541.3 watts x 3 hours x 30 days = 1218.171 kWh used in a month now multiply price per kwh, 8.75 * 1218.171 = 10658.996 philipine pesos for the monthly bill

OpenStudy (anonymous):

you can save a lot of money by not using airconditioner

OpenStudy (anonymous):

@IrishBoy123 is this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the sum of all add is incorrect, our professor told to use the 1hp =746 watts @brett.dodd16

OpenStudy (anonymous):

well u might be reading this all wrong

OpenStudy (anonymous):

use 1492 then

OpenStudy (anonymous):

for what

OpenStudy (anonymous):

"but the sum of all add is incorrect, our professor told to use the 1hp =746 watts @brett.dodd16 "

OpenStudy (anonymous):

2hp = 2*276 watts = 1492 watts